Find the minimal polynomial of T: R^3→R^3 defined by T(a,b,c)= (a-b,b,c)


1
Expert's answer
2021-02-24T07:01:00-0500

We seek eigenvalues of this operator:

"T(a,b,c)=(a-b,b,c)=(\lambda a, \lambda b, \lambda c)".

This implies that "(\lambda-1)b=(\lambda-1)c=0" and "(\lambda-1)a=-b".

The only value of "\lambda" which provides a non-zero solution is "\lambda=1".

Therefore, the characteristic polynomial of T is (x-1)3.

The minimal polynomial is a factor of the characteristic polynomial (x-1)3.

It may be (x-1)3, (x-1)2 or x-1. Check them all.

"(T-E)(a,b,c)=(-b,0,0)", T-E is not a zero operator.

"(T-E)^2(a,b,c)=(T-E)(-b,0,0)=(0,0,0)", hence, (T-E)2=0.

Therefore, the minimal polynomial of T is (x-1)2.